Shift report 20011104 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Nov 04 2001 - 00:08:29 EST

  • Next message: Apache: "Shift report 20011104 00:00-08:00"

    ------------------------------------------------------------
    Date: Sun Nov  4 00:08:29 2001, Shift: 16:00-24:00
    Supervisor: sjs, with: tst
    Log book#10 pages 92 - 92
    
    Shift Summary:
    ------------------------------------------------------------
    1.) Finally good news--machine running well throughout 
    shift.
    2.) Data run 5649 (trig 1/6 = 280k/152k),
             run 5650 (trig 1/6 = 277k/85k)
             run 5654(started)
        Original store at 16:30 started with ZDC rate around 
        850, rapidly dropping to under 400-500, then leveling
        off.   New store at 23:30 started with ZDC rate of
        830, down to 600 at midnight...
    3.) No new problem uncovered.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Nov 04 2001 - 00:09:15 EST